# Hydra: Virtualized Multi-Language Runtime for High-Density Serverless Platforms

### v2

@@ Abstract.
Serverless is an attractive computing model that offers seamless scalability and elasticity; it takes the infrastructure management burden away from users and enables a pay-as-you-use billing model. As a result, serverless is becoming increasingly popular to support highly elastic and bursty workloads. However, existing platforms are supported by bloated virtualization stacks which, combined with bursty and irregular invocations, leads to high memory and latency overheads.
To reduce the virtualization stack bloat, we propose Hydra, a virtualized multi-language serverless runtime capable of handling multiple invocations of functions written in different languages. To measure its impact in large platforms, we build a serverless platform that optimizes scheduling decisions to take advantage of Hydra by consolidating function invocations on a single instance, reducing the total infrastructure tax. Hydra improves the overall function density (ops/GB-sec) by 4.47$\times$ on average compared NodeJS, JVM, and CPython, the state-of-art single-language runtimes used in most serverless platforms. When reproducing the Azure Functions trace, Hydra reduces the overall memory footprint by 2.1 $\times$ and reduces the number of cold starts between 4 and 48 $\times$.
@@ 1. Introduction
Figure 1. Startup latency and memory footprint of different virtualization systems (Refer to caption), runtimes (Refer to caption), the impact of forking and snapshotting (Refer to caption), and Hydra Sandbox (Refer to caption).
1. Serverless runs on bloated virtualization stacks. A container (e.g., Docker (docker, )) or a Virtual Machine (e.g., QEMU (qemu, ), Firecracker (agache:2020, )), and a language runtime (e.g., NodeJS, CPython, or JVM) running on top of it is the most commonly used virtualization stack. For each invocation being handled by the platform, a separate virtualization stack is required. State-of-the-art container/VM implementations require a long startup time and a high memory footprint compared to the average mean function invocation latency and memory footprint (sharad:2020, ) (see Figure 1). The same applies to language runtimes such as CPython, JVM, and NodeJS, the three most popular native language implementations that, according to a recent study (newrelic, ), account for approximately 90% of the total number of function invocations in Amazon Lambda (lambda, ). These runtimes were developed to host long-running applications and, therefore, employ effective but time and resource-consuming techniques to optimize for long-term performance (carreira:2021, ). The combined memory footprint and startup latency of container/VM and runtime can easily dominate the total memory footprint and startup latency for lightweight serverless functions.
2. Serverless workloads are bursty. Such workloads include periods of very low activity with only a few invocations, followed by periods of higher activity with multiple function invocations (sharad:2020, ). Figure 2 shows how sparse the invocations of each function are; only 11% of the functions will receive on average at least 1 invocation per minute, meaning that most virtualization stacks are highly under-utilized. As a result of these factors, platforms resort to a tradeoff between high memory overheads to cache virtualization stacks (waiting for the next invocation), and long tail invocation latencies inflicted by the cost of launching a new virtualization stack (cold start). Figure 3 depicts the effect of caching functions that have been executed recently. Due to the sparsity of invocations of each function, the number of cached functions is much higher than the number of invocations.
@@
Checkpoint/Restore (shin:2022, ; ustiugov:2021, ; silva:2020, ) (C/R) and runtime forking (akkus:2018, ; oakes:2018, ) have been extensively studied as a way to reduce memory footprint and tail latency. However, while both C/R and runtime forking reduce memory and startup latency by sharing a common stack (see Figure 1), both techniques only optimize invocations of a single function requiring separate virtualization stacks when multiple functions are invoked. Other techniques supporting lightweight sandboxes (dukic:2020, ) still only allow single function invocations to co-execute on the same runtime, therefore requiring separate virtualization stacks for different functions.
We propose Hydra  [^1] Hydra is a Greek mythology creature with multiple heads that we use as an analogy for the multi-function and multi-language support., a runtime capable of sharing a single virtualization stack across concurrent invocations of multiple functions, even if written in different languages. Hydra supports both AOT-compiled and JIT-compiled functions, and function invocations are executed in a lightweight sandbox that can be snapshotted and restored. Opposed to previous VM or process C/R that snapshot the entire system or process, Hydra supports C/R of a single sandbox, allowing cached state and compiled code of a single function to be restored at any point.
On a number of established serverless function benchmarks (dukic:2020, ; copik:2021, ), we demonstrate how Hydra’s lightweight sandboxes increase function density (ops/GC-sec) in single-function scenarios compared to VM C/R, process forking, and running functions in separate stacks. To study the impact of Hydra on real-world platforms, we built a platform that schedules invocations of each tenant to the same Hydra instance. We show that a tenant-oriented scheduling algorithm significantly reduces overall memory consumption by 2.1 $\times$ and reduces the number of cold starts by 4/48 $\times$ when reproducing a public serverless trace (sharad:2020, ).
@@
In summary, our contributions are the following:
  • we demonstrate that serverless cache utilization can be improved by re-using the same virtualization stack for different functions (section 1);
  • we build the first (to the best of our knowledge) serverless runtime capable of sharing a single virtualization stack across multiple functions written in different languages while offering C/R support for individual sandboxes (section 3);
  • we build a serverless platform that incorporates a tenant-aware scheduling algorithm to show how Hydra could be deployed at scale (section 4);
  • we measure the performance impact of Hydra both at a micro-scale by comparing its density with single-language runtimes, and also at a macro-scale by measuring its impact while replaying the Azure Functions trace on a realistic platform (section 5).
2. Background
Serverless platforms inherit much of the virtualization techniques that were initially designed to support long-running monolithic applications and microservices. We argue that this infrastructure is not a good fit for lightweight, short-running, and bursty serverless functions.
@@ 2.1. Traditional Virtualization Stacks
Snapshotting and forking have also received much attention as a promising technique to reduce the memory footprint and startup latency of serverless functions. Restoring from a previous snapshot significantly reduces the startup time and memory footprint compared to a cold start (du:2020, ; ustiugov:2021, ; cadden:2020, ) but still requires a separate virtualization stack for each concurrent invocation and imposes storage/network overhead to manage snapshots (ao:2022, ). Similarly, runtime forking (oakes:2018, ; akkus:2018, ) also reduces both memory footprint via copy-on-write and startup time by simply forking instead of launching a completely new virtualization stack. However, state-of-the-art runtimes such as CPython, NodeJS, and JVM do not support forking out-of-the-box. For example, Garbage Collector and Just-In-Time compiler threads do not survive the forkin and require careful revival in the child process.
@@ 2.2. Lightweight Execution Environment
Allowing multiple functions to run in a single runtime reduces the startup latency and memory footprint of serverless functions by using a single virtualization stack. However, it requires enforcing resource isolation. CPU, file system, and network isolation can be guaranteed using Operating System-level primitives such as Linux Namespaces and Control Groups (cgroups, ). Memory isolation requires function sandboxing, which can then be further strengthened with hardware features such as MPK  (mpk, ; libmpk, ; erim, ; kotni:2021, ) and CHERI (woodruff:2014, ; vasily:2022, ). In this paper, we focus on how to sandbox functions efficiently.
@@
There are two main widely available sandboxing techniques: WebAssembly and Memory Isolates. WebAssembly (haas:2017, ) is a binary-code format with memory safety guarantees that resulted from the evolution of previous Software Fault Isolation techniques such as NaCL (nacl, ). These guarantees are enforced by constraining memory access to a single linear byte array with efficient bounds checks during compilation and run-time. Runtimes supporting the execution of WebAssembly functions could use it to run multiple functions in a memory-isolated environment (shillaker:2020, ). However, for functions that are interpreted (such as Python), the function code would still have to include an interpreter (CPython, for example), re-introducing resource redundancy as invocations do not share the same interpreter (shillaker:2020, ).
@@
Language runtimes, such as V8 (v8, ) and GraalVM Native Image (graalvm, ), offer Memory Isolates, a memory segment that can be used to host the execution of a function. Similarly to WebAssembly sandboxes, an isolate is a linear memory segment used to keep application objects. References are resolved within the linear memory segment and the sandbox limits are enforced by the compiler and runtime. Isolates are, however, restricted to executing functions written in the language supported by the runtime, JavaScript for V8, and Java for GraalVM Native Image.
@@ 2.3. Multi-Language Support
Executing functions developed in different languages on top of a single runtime instance is also a key ingredient to reducing virtualization stack redundancy. WebAssembly accomplishes this by allowing multiple languages to target its binary-code format. On the other hand, memory isolates alone cannot offer multi-language capabilities on itself. An alternative approach is enabled by using Truffle (wuerthinger:2013, ), a language implementation framework that allows developers to quickly write interpreters for a specific language. Truffle interpreters are then automatically processed and optimized, taking full advantage of Truffle’s underlying JIT compiler. Truffle is supported by recent advances in language runtime and compiler literature (wimmer:2019, ; grimmer:2015a, ; latifi:2021, ; grimmer:2015b, ; zhang:2014, ; grimmer:2014, ; wuerthinger:2013, ; wimmer:2012, ; wuerthinger:2017, ; wuerthinger:2013, ; larisch:2018, ) and is designed to efficiently execute dynamic languages such as Python, JavaScript, and Ruby but also has support for LLVM bitcode (rigger:2016, ) and even WebAssembly (salim:2020, ). Truffle allows functions written in different languages to execute in a Java runtime. However, Truffle code execution greatly suffers from long initialization times and high memory overhead (section 5.1.3), which easily nullifies any benefit obtained by virtualizing the runtime.
3. Virtualized Multi-Language Runtime
Hydra is a serverless runtime capable of running multiple functions in isolation by confining each function inside a lightweight sandbox. Figure 4 depicts a Hydra instance with 3 active sandboxes of 2 different functions. Each sandbox runs invocations of a single function, one invocation at a time. At startup, Hydra opens a network socket to serve REST requests to register, deregister, and invoke functions (section 3.1). Upon receiving an invocation request, Hydra loads the function into memory if not already loaded, creates a new sandbox if no warm sandbox is available, and invokes the function inside the sandbox. Sandboxes of the same function share the same code cache and the initial heap state (section 3.2). Hydra introduces a function framework that automatically uses Truffle (section 3.3), and supports launching new sandboxes or restoring one from a previous checkpoint (section 3.4).
Figure 4. Hydra running 3 sandboxes (Memory Isolates).
@@ 3.1. Function Interface
Hydra exposes a minimal canonical interface to the outside through a single end-point. Only three methods are required, function registration, invocation, and deregistation:
{minted}
[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, ]java boolean regFunction(byte[] code, String functionID, String functionEntrypoint); String invkFunction(String funID, String arguments); boolean delFunction(String fuctionID);
{listing*}
[!ht] {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, ]c
int create_isolate(create_isolate_params_t* params, isolate_t** isolate); int attach_thread(isolate_t* isolate, isolatethread_t** thread); int copy(isolate_t* isolate, char* buffer, long size, handle_t** handle); int invoke_function(isolate* guest, isolate* host, handle_t* entrypoint, handle_t* arguments, handle_t** result); int detach_thread(isolate_t* isolate, isolatethread_t* thread); int tear_down_isolate(isolate_t* isolate);

### v3

@@ Abstract.
Serverless is an attractive computing model that offers seamless scalability and elasticity; it takes the infrastructure management burden away from users and enables a pay-as-you-use billing model. As a result, serverless is becoming increasingly popular to support highly elastic and bursty workloads. However, existing platforms are supported by bloated virtualization stacks, which, combined with bursty and irregular invocations, lead to high memory and latency overheads.
To reduce the virtualization stack bloat, we propose Hydra, a virtualized multi-language runtime and platform capable of hosting multiple sandboxes running concurrently. To fully leverage Hydra’s virtualized runtime, we revisit the existing serverless platform design to make it colocation-aware across owners and functions, and to feature a caching layer of pre-allocated Hydra instances that can be used by different functions written in different languages to reduce cold starts. We also propose a snapshotting mechanism to checkpoint and restore individual sandboxes.
[…]
@@ 1. Introduction
Figure 1. Performance landscape of different virtualization systems (green), language runtimes (blue), impact of forking and snapshotting (red), and Hydra Sandbox (black).
1. Serverless runs on bloated virtualization stacks. A container (e.g., Docker (docker, )) or a Virtual Machine (e.g., QEMU (qemu, ), Firecracker (agache:2020, )), and a language runtime (e.g., Node.js, CPython, or JVM) running on top of it are the most commonly used virtualization stacks. For each invocation handled by the platform, a separate virtualization stack is provisioned. State-of-the-art container/VM implementations lead to a long startup time and a high memory footprint inflicting a non-negligible overhead on lightweight and fast-execution serverless functions. The same applies to language runtimes such as CPython, JVM, and Node.js, the three most popular native language implementations that, according to a recent study (newrelic, ), account for approximately 90% of the total number of function invocations in Amazon Lambda (lambda, ). These runtimes were developed to host long-running applications and, therefore, employ effective but time and resource-consuming techniques to optimize for long-term performance (carreira:2021, ; kohli:2024, ). The combined memory footprint and startup latency of a container/VM and a runtime can easily dominate the total memory footprint and startup latency for serverless functions, that have been consistently reported as lightweight and fast executing in public traces (sharad:2020, ; joosen:2023, ; wang:2021, ).
2. Serverless workloads are bursty. Serverless workloads include periods of very low activity with only a few invocations, followed by periods of higher activity with multiple function invocations (sharad:2020, ). Figure 2 shows how sparse the invocations of each function are; only 11% of the functions will receive on average at least 1 invocation per minute, meaning that most virtualization stacks are highly underutilized. As a result of these factors, platforms resort to a trade-off between high memory overheads to cache virtualization stacks (waiting for the next invocation) and long tail invocation latencies inflicted by the cost of launching a new virtualization stack (cold start). Figure 3 depicts the effect of caching functions recently invoked, a commonly used technique in both production systems such as Amazon Lambda, Google Cloud Functions, and Azure Functions, and open source serverless platforms such as OpenWhisk and Knative (wang:2018, ; lin:2025, ). The results demonstrate that, as a consequence of the sparsity of the invocations of each function, the number of cached functions is much higher than the number of invocations. In turn, platforms need to carefully balance the size of the cache (memory overhead) with the number of function cold starts (tail latency).
@@
Checkpoint/Restore (shin:2022, ; ustiugov:2021, ; silva:2020, ) (C/R) and runtime forking (akkus:2018, ; oakes:2018, ) have been extensively studied as a way to reduce memory footprint and tail latency. However, while both C/R and runtime forking reduce memory and startup latency by sharing a common stack (see Figure 1), both techniques optimize invocations of a single function, requiring separate virtualization stacks when multiple functions are invoked. Other techniques already virtualize language runtime (dukic:2020, ) but only allow single function invocations to co-execute on the same runtime, therefore requiring separate virtualization stacks for different functions.
In this paper, we propose Hydra, a runtime and platform capable of sharing a single virtualization stack across concurrent invocations of multiple functions, even if written in different languages. Hydra’s runtime supports both AOT-compiled and JIT-compiled functions, and function invocations are executed in lightweight sandboxes that can be checkpointed and restored individually. In addition, we propose revisiting the serverless platform design to make it colocation-aware, allowing it to be configured with different colocation policies such as colocating all functions from the same tenant, or only colocating invocations for the same function.
[…]
@@
In summary, this work offers the following contributions:
  • we demonstrate that density and elasticity can be significantly improved with Hydra. Density is improved by sharing additional components of the virtualization stack (section 3). Elasticity is improved restoring individual sandboxes from individual sandbox snapshots;
  • we rethink the serverless platform design to incorporate colocation-aware scheduling and elastic runtime pools that effectively mitigate cold starts;
  • we measure the elasticity and density benefits of Hydra both at a local scale by comparing it to state-of-the-art language runtimes, and at a distributed scale by replaying the Azure Functions trace (sharad:2020, ) on a realistic platform (section 5).
2. Background and Motivation
Serverless platforms inherit much of the virtualization technology that was designed to support long-running monolithic applications and microservices. In addition, serverless platforms have also traditionally been designed to work with virtualization stacks that do not support colocation. We argue that this infrastructure is not a good fit for lightweight, short-running, and bursty serverless functions.
@@ 2.1. Traditional Virtualization Stacks
Snapshotting and forking have also received much attention as a promising technique to reduce the memory footprint and startup latency of serverless functions. Restoring from a previous snapshot significantly reduces the startup time and memory footprint compared to a cold start (du:2020, ; ustiugov:2021, ; cadden:2020, ) but still requires a separate virtualization stack for each concurrent invocation and imposes storage/network overhead to manage snapshots (ao:2022, ). Similarly, runtime forking (oakes:2018, ; akkus:2018, ) also reduces both memory footprint via copy-on-write and startup time by simply forking instead of launching a completely new virtualization stack. However, state-of-the-art runtimes such as CPython, Node.js, and JVM do not support forking out-of-the-box. For example, Garbage Collector and Just-In-Time compiler threads do not survive forking and require careful revival in the child process.
@@ 2.2. Lightweight Execution Environment
Running multiple function invocations in a single runtime reduces startup latency and memory footprint of serverless functions by using a single virtualization stack. However, it requires enforcing resource isolation. CPU, file system, and network isolation can be guaranteed using Operating System-level primitives such as Linux Namespaces and Control Groups (cgroups, ). Memory isolation requires function sandboxing.
@@
There are two main widely available sandboxing techniques in modern language engines: WebAssembly sandboxes and Memory Isolates. WebAssembly (haas:2017, ) is a binary-code format with memory safety guarantees that resulted from the evolution of previous Software Fault Isolation techniques such as NaCL (nacl, ). These guarantees are enforced by restricting memory access to a single linear byte array with efficient bound checks during compilation and run-time. Runtimes supporting the execution of WebAssembly sandboxes can run multiple functions in a memory-isolated environment (shillaker:2020, ; gadepalli:2020, ). However, modern WebAssembly runtimes do not include high-level languague support and thus, the function code would have to include the language engine. For example, existing support to run Python/JavaScript code requires functions to include CPython/V8 (or other engine) compiled into WebAssembly (shillaker:2020, ; fermyon, ). This mechanism re-introduces redundancy as sandboxes do not share language engines, making it similar to OS-level virtualization stacks in Figure 4.
@@
Language runtimes such as V8 (v8, ) and GraalVM Native Image (graalvm, ) offer Memory Isolates, a memory segment that can be used to host the execution of a function. Similarly to WebAssembly sandboxes, an isolate is a linear memory segment used to keep application objects. References are resolved within the linear memory segment and the sandbox limits are enforced by the compiler and runtime. Isolates are, however, restricted to executing functions written in the language supported by the runtime, JavaScript for V8, and Java bytecode for GraalVM Native Image.
Platforms such as Photons (dukic:2020, ) and Knative (knative, ) propose weakening isolation to run concurrent requests in a single runtime with minimal or even no isolation at all. Photons offers transparent isolation for Java applications by applying Java bytecode instrumentation to automatically synchronize access to global shared state. Knative, on the other hand, leaves this concern to the developers, forcing serverless application developers to deal with concurrency and synchronization challenges.
Figure 4. Virtualization stacks (gray boxes) needed to serve two concurrent invocations of two functions in different serverless platforms.
2.3. Runtime Impact on Serverless Platforms
The execution environment used to run functions can affect the infrastructure. For example, as shown in Figure 4, in order to run two concurrent invocations to two different functions, the serverless platform has to create four OpenWhisk (openwhisk, ) workers since OpenWhisk runtimes isolate each concurrent invocation in a single worker. If the platform uses Knative workers, it needs only two workers since Knative is capable of colocating concurrent invocations to the same function, providing no isolation within the same worker. Hydra, on the other hand, requires a single worker to execute all invocations of both functions.
Below, we illustrate how the type of runtime can affect the serverless design of the serverless platform:
1. Limiting function resources. Serverless workers running different functions share the same hardware. Therefore, it is crucial to limit worker resources, such as memory, to ensure a fair distribution of hardware between functions. If a serverless worker runs one request at a time, then its memory can be restricted to a fixed value. However, if the worker supports handling concurrent invocations (as in Photons or Knative), its memory limits have to grow and shrink to accommodate function invocations bursts. However, modern language runtimes such as the JVM, and NodeJS do not support dynamic memory heap resizing, leading to significant memory waste (bruno:2018, ).
2. Pre-creating serverless workers. Platforms like OpenWhisk feature function-generic language-specific container images for workers. A function is registered in the available worker before handling invocation. This enables the platform to pre-create a language-partitioned pool of workers to avoid creating a VM/container on the critical path of a cold start. However, this requires the platform to resize the worker’s memory to the function’s needs after creation. Language runtimes such as JVM probe the total available memory during startup to determine their heap size parameters, and reconfiguring the heap size at run-time is not trivial (bruno:2018, ), preventing platforms from effectively resizing the worker. Other platforms, such as Knative, require developers to provide function-specific container images to start workers. This approach prevents platforms from keeping a pool of pre-created generic workers.
@@ 2.4. Multi-Language Support
Executing functions developed in different languages on top of a single runtime instance is also a key ingredient to reducing virtualization stack redundancy. WebAssembly accomplishes this by allowing multiple languages to target its binary code format or by including interpreters compiled into WebAssembly in the function code. On the other hand, memory isolates alone cannot offer multi-language capabilities on itself. An alternative approach is enabled by using Truffle (wuerthinger:2013, ), a language implementation framework that allows developers to quickly write interpreters for a specific language. Truffle interpreters are then automatically processed and optimized, taking full advantage of Truffle’s underlying JIT compiler. Truffle is supported by recent advances in language runtime and compiler literature (wimmer:2019, ; grimmer:2015a, ; latifi:2021, ; grimmer:2015b, ; zhang:2014, ; grimmer:2014, ; wuerthinger:2013, ; wimmer:2012, ; wuerthinger:2017, ; wuerthinger:2013, ; larisch:2018, ) and is designed to efficiently execute dynamic languages such as Python, JavaScript, and Ruby but also has support for LLVM bitcode (rigger:2016, ) and even WebAssembly (salim:2020, ). Truffle allows functions written in different languages to execute in a Java runtime. However, Truffle code execution greatly suffers from long initialization times and high memory overhead (section 5.1.3), which easily nullifies any benefit obtained by virtualizing the runtime.
3. Hydra Language Runtime
Hydra is a language runtime capable of running multiple functions concurrently and in isolation, by confining each function into a lightweight sandbox. Figure 5 depicts an example of a Hydra instance with 3 active sandboxes executing 2 different functions. Each sandbox runs invocations of a single function, one invocation at a time. At startup, Hydra opens a network socket to serve REST requests (section 3.1). Upon receiving an invocation request, Hydra loads the function into memory if not already loaded, creates a new sandbox if no warm sandbox is available, and invokes the function inside the sandbox. Sandboxes of the same function share the same code cache and initial heap state (section 3.2). Hydra offers multi-language support through Truffle (section 3.3), and supports launching new sandboxes or restoring one from a previous snapshot (section 3.4).
Figure 5. Hydra running three function sandboxes.
@@ 3.1. Function Interface
Hydra exposes a minimal canonical interface to the outside through a single end-point. Only three methods are required, function registration, invocation, and deregistation. This function interface is heavily inspired by existing function interfaces available in both commercial (Amazon Lambda (lambda, ), Azure Functions (functions, ), and Google Cloud Functions (gcf, )), and open-source serverless platforms (such as OpenFaaS (openfaas, ), OpenWhisk (openwhisk, ), Knative (knative, )). We expect that Hydra requires minor to no function source code modifications to run existing serverless functions in Hydra.
3.2. Function Sandboxes
Hydra sandboxes are based on GraalVM’s Native Image (wimmer:2019, ) memory isolates which we extend to support on-demand sandbox creation, function invocation, and checkpoint/restore (section 3.4). Native Image is an Ahead-Of-Time Java compiler that we use to compile user functions into shared libraries. Hydra’s design is independent from the sandbox implementation and could have been based on WebAssembly sandboxes, for example.
Upon receiving a request to invoke a function, Hydra loads the function shared library using dlopen, if not already loaded. The shared library loads into memory the function code cache and the initial heap state (see Figure 5). The code cache contains the compiled user code that will be invoked through the registered entrypoint and the initial heap state contains a set of global values which are copied into the sandboxe’s memory heap upon modification (wimmer:2019, ) (copy-on-write). Both the code cache and the initial heap state are shared among all sandboxes of a single function. Different functions have separate code caches and initial heap states.
Listing 3.2 presents the ABI that Hydra uses to create and invoke function sandboxes.  [^1] Hydra’s ABI is built on top of Native Image shared library ABI. The create_sandbox creates a sandbox and returns a pointer to the new sandbox (through the second function argument). Once a sandbox is created, threads can be attached to it through the attach_thread function. This function creates a new thread stack in the given sandbox and returns a pointer to the thread data structure (which can be later used to detach the thread from the sandbox by calling detach_thread). Note that threads can be created in any sandbox and will start executing in the same sandbox as the parent thread.
{listing}
[!ht] {Verbatim}[commandchars=
{}] int create˙sandbox(sandbox˙params˙t*, sandbox˙t**); int attach˙thread(sandbox˙t*, sandbox˙thread˙t**); int invoke˙sandbox(sandbox˙thread˙t*, char*, char*, int); int detach˙thread(sandbox˙thread˙t*); int tear˙down˙isolate(sandbox˙t*);